### v1

# A Simple Approach to Constraint-Aware Imitation Learning with Application to Autonomous Racing

Abstract
Guaranteeing constraint satisfaction is challenging in imitation learning (IL), particularly in tasks that require operating near a system’s handling limits. Traditional IL methods often struggle to enforce constraints, leading to suboptimal performance in high-precision tasks. In this paper, we present a simple approach to incorporating safety into the IL objective. Through simulations, we empirically validate our approach on an autonomous racing task with both full-state and image feedback, demonstrating improved constraint satisfaction and greater consistency in task performance compared to a baseline method.
I Introduction
Autonomous racing competitions, including the Indy Autonomous Challenge, F1TENTH, and the DARPA Grand Challenge, drive innovation in high-speed control and real-time decision-making [av-racing-survey]. Traditional control strategies focus on optimal race-line generation but require extensive engineering and prior knowledge. Reinforcement learning (RL) can outperform human drivers in drone [kaufmann_rl_drone] and simulated vehicle racing [nature-rl-car], yet demands risky, large-scale training and struggles with sim-to-real transfer. Moreover, state-of-the-art platforms often rely on advanced sensors (e.g., LiDAR, D-GPS) and significant onboard computation, limiting deployment on low-cost systems. Consequently, developing efficient, learning-based control for resource-constrained hardware and limited sensing remains challenging.
Imitation Learning (IL), which has demonstrated success in robotics, autonomous vehicles, and gaming [imitation-learning-survey-robotics, imitation-learning-survey-av, imitation-learning-survey-gaming], emerges as a promising candidate to address these constraints. Recent advances in deep learning have further enhanced IL performance, allowing agents to mimic complex human strategies through the use of high-capacity models that map raw sensor data (e.g., images) to actions [pan_end2end_il, loquercio_deep_drone_acrobatics].
Despite these successes, IL-based methods often overlook safety-critical aspects of decision-making. In many real-world scenarios, such as autonomous driving and drone navigation, the consequences of safety constraint violations can be severe. Purely maximizing imitation accuracy can lead to suboptimal or unsafe maneuvers, especially if the agent model has insufficient representation power to perfectly clone the expert demonstrations in all possible states. Consequently, ensuring that an IL agent respects safety constraints (e.g., collision avoidance, maintaining certain system limits) remains an open and pressing challenge.
To address safety concerns, several recent efforts have attempted to augment IL with constraint satisfaction mechanisms. Many of these methods assume that the expert demonstrations themselves are safe or reliable, thereby focusing on identifying and prioritizing safety-critical states in the dataset [safe_gil, spencer_covariate_shift]. By emphasizing these states, the agent’s learned policy can allocate higher importance to correct behavior under dangerous conditions. SafeDAgger [safe-dagger] takes a different approach by training an additional classifier to predict deviations from the expert and switching to a backup policy when necessary. However, these approaches become less effective if the expert demonstrations are themselves risky or even contain unsafe actions. This gap highlights the need for a more robust framework that can handle unsafe expert trajectories while still aiming to learn a safe policy.
Parallel research in classical control has long studied the problem of safe control through approaches such as Model Predictive Control (MPC), Control Barrier Functions (CBFs), and Hamilton-Jacobi reachability analysis [ugo_lmpc, reference-governor, bansal_hj_reach, ad_cbf]. These methods provide rigorous guarantees about constraint satisfaction by, for example, constructing safe sets or using invariant sets to ensure that the system remains within specified limits. A common strategy involves the use of safety filters or shielding layers that modify the control commands if a safety violation is imminent [predictive_safety_filter]. Although such formal methods offer strong theoretical guarantees, they often scale poorly to high-dimensional, nonlinear systems. The computational overhead of solving online optimization problems or explicitly computing reachable sets becomes prohibitive in complex environments.
In this work, we address the problem of constraint-aware imitation learning under the realistic assumption that expert demonstrations may be unreliable or partially unsafe. Specifically, we aim to (1) learn a policy that mimics the behaviors from a set of expert demonstrations (some of which are failures) while avoiding constraint violation, and (2) develop a safety mechanism in the training pipeline that can handle image-based inputs and partial state observations.
Our main contributions are as follows:
  • Novel Implicit Safety Filter Architecture: We propose a differentiable safety filter that can be embedded into classical imitation learning frameworks for high-dimensional, partially observed systems.
  • Data-driven Approximation of Safe Set: We introduce a simple data-driven approach to approximate the safe set for constrained systems inspired by [ugo_lmpc].
  • Empirical Validation: We demonstrate the effectiveness of our approach on high-dimensional, nonlinear, image-feedback systems, highlighting improvements in constraint satisfaction and consistency in task performance compared to baseline methods.
II Problem Formulation
II-A System
Consider a nonlinear, time-invariant, deterministic, discrete-time system described as:
$x_{k+1}=f(x_{k},u_{k}),$ (1)
$y_{k}=h(x_{k})+n_{k},$
where $x_{k}$ is the state, $u_{k}$ is the control input, $y_{k}$ is the measurement, and $n_{k}$ is the measurement noise at time step $k$. The system is subject to constraints defined as:
$x_{k}\in\mathcal{X},~{}u_{k}\in\mathcal{U},~{}\forall k\geq 0.$ (2)
As a shorthand notation, we denote $\mathbf{u}_{0:N-1}=\{u_{0},\dots,u_{N-1}\}$ as a control sequence. The state at time $N$ resulting from applying the control sequence $\mathbf{u}_{0:N-1}$ to the system with dynamics $f$, initialized at the state $x_{0}$, is denoted by $x_{N}=f(x_{0},\mathbf{u}_{0:N-1})$, where each intermediate state is determined recursively with (1). We also denote a closed-loop trajecotry by $\mathbf{x}_{0:N}=\{x_{0},\dots,x_{N}\}$. This implies that there exists a feasible control sequence $\mathbf{u}_{0:N-1}$ such that $x_{k+1}=f(x_{k},u_{k})$ for all $k=0,\dots,N-1$.
II-B Constrained Optimal Control Task
In this paper, we consider the following constrained optimal control problem.
$\min_{\mathbf{u}_{0:\infty}}{}{}$ $\sum_{k=0}^{\infty}c(x_{k},u_{k})$ (3)
s.t., $x_{k+1}=f(x_{k},u_{k}),$
$y_{k}=h(x_{k})+n_{k},$
$x_{k}\in\mathcal{X},~{}u_{k}\in\mathcal{U},$
$\forall k\geq 0.$
where $c(\cdot,\cdot)$ is a cost function. We assume the existence of a zero-cost target set $\mathcal{X}_{f}$: once the system reaches $\mathcal{X}_{f}$, it can remain in the set at no further cost. Specifically, we use the following assumption on $c(\cdot,\cdot)$ and $\mathcal{X}_{f}$.
Assumption 1
(Non-negative Cost Function and Target Set) The cost function $c(x,u)$ is non-negative for all $x\in\mathcal{X},u\in\mathcal{U}$. Moreover, $\forall x\in\mathcal{X}_{f}$, $\exists u\in\mathcal{U}$, s.t., $f(x,u)\in\mathcal{X}_{f}$, $c(x,u)=0$.
We determine whether a trajectory is successful by checking whether its terminal state reaches the target set, namely,
Definition 1
A closed-loop trajectory $\mathbf{x}_{0:N}=\{x_{0},\dots,x_{N}\}$ is called successful if $x_{N}\in\mathcal{X}_{f}$, and failed if $x_{N}\notin\mathcal{X}_{f}$.
II-C Safe Imitation Learning Problem
We are focused on the case where the controller only has access to the output $y_{k}$ rather than direct access to $x_{k}$. Our objective is to design an output-feedback controller $\pi_{\theta}(y)$ that approximately solve (3). Assume we have access to a high-performing full-state feedback policy, denoted as $\pi_{\beta}(x)$, which acts as the expert. Our approach is to mimic $\pi_{\beta}(x)$ for designing the output feedback controller $\pi_{\theta}(y)$. Formally, this objective can be expressed as:
$\min_{\theta}{}{}$ $\mathbb{E}_{(x,y)\sim P((x,y)\mid\theta)}\left[\mathcal{L}(\pi_{\theta}(y),\pi_{\beta}(x))\right],$ (4)
s.t. $f(x,\pi_{\theta}(y))\in\mathcal{R}_{\infty}^{B}(\mathcal{X}_{f}),$
where $P((x,y)|\theta)$ denotes the distribution of state-output pairs induced by $\pi_{\theta}$, $\mathcal{L}(\cdot,\cdot)$ is a measurement of the discrepancy between two actions, and $\mathcal{R}_{\infty}^{B}(\mathcal{X}_{f})$ is a backward reachable tube from the target set $\mathcal{X}_{f}$, which is defined as follows.